Find the Laplace transform of y"-4y'+9y=t ,y(0)=0,y'(0)=1
1
Expert's answer
2020-05-24T18:55:24-0400

Laplace transform of  y"-4y'+9y=t , is given as

L( y"-4y'+9y=t)=L(y'')-L(4y')+L(9y)=L(t)

L(y'')="s^2F(s)-sf(0)-f'(0)"

It is given in the question that, y(0)=0 and y'(0)=1.

L(y)="Y(s)"

Hence, L(y'') = "s^2Y(s)-s.0-1=s^2Y(s)-1."

Similarly, L(y')="sY(s)-y(0)=sY(s)"

L(t)="\frac{1}{s^2}"

Putting the above values in the differential equations,

we get,

"s^2Y(s)-1-4sY(s)+9Y(s)=\frac{1}{s^2}"

or, "Y(s) (s^2-4s+9)=\frac{1}{s^2}+1"

or, "Y(s)=\frac{1+s^2}{s^2(s^2-4s+9)}" (Answer)